# Total Gluon Helicity Contribution to Proton Spin from Lattice QCD

### v1

@@ Acknowledgment
$\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}\equiv\begin{pmatrix}\bar{Z}^{\mathrm{DR}}_{11}&\bar{Z}^{\mathrm{DR}}_{12}\\ \bar{Z}^{\mathrm{DR}}_{21}&\bar{Z}^{\mathrm{DR}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,DR}}_{11}&Z^{\mathrm{RI,\,DR}}_{12}\\ Z^{\mathrm{RI,\,DR}}_{21}&Z^{\mathrm{RI,\,DR}}_{22}\end{pmatrix}^{-1}\Big|_{\epsilon\rightarrow 0}=\begin{pmatrix}\bar{Z}^{\mathrm{lat}}_{11}&\bar{Z}^{\mathrm{lat}}_{12}\\ \bar{Z}^{\mathrm{lat}}_{21}&\bar{Z}^{\mathrm{lat}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,lat}}_{11}&Z^{\mathrm{RI,\,lat}}_{12}\\ Z^{\mathrm{RI,\,lat}}_{21}&Z^{\mathrm{RI,\,lat}}_{22}\end{pmatrix}^{-1}\,.$ (27)
Consequently, the $R_{ij}$ determined using the first equality $\bar{Z}^{\mathrm{DR}}\cdot\big(Z^{\mathrm{RI,\,DR}}\big)^{-1}\big|_{\epsilon\rightarrow 0}$ with DR can be employed, together with the input $Z^{\mathrm{RI,\,lat}}_{ij}\equiv Z^{\mathrm{RI}}_{ij}$ computed in (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD), to obtain the values for the $\overline{\mathrm{MS}}$-renormalization constants $\bar{Z}^{\mathrm{lat}}_{ij}$ of the operator $K^{\mu}$ and $J_{5}^{\mu}$ but with UV-divergences regularized using Lattice, which can then be applied directly to the bare operator matrix elements computed in Lattice to obtain the $\overline{\mathrm{MS}}$-renormalized results. By virtue of (3), we can further derive the following shortcut expressions for $R_{ij}$:
@@
where $\overline{K}(\overline{J})_{gg(\bar{q}q)}$ denote the Lorentz-invariant light-cone form-factors or matrix elements of the $\overline{\mathrm{MS}}$-renormalized $K(J)$-current operator but with $\mathrm{RI/MOM}$-renormalized external $g(q)$-fields. Note that $K^{\mathrm{RI}}_{gg}$ and $J^{\mathrm{RI}}_{\bar{q}q}$ equal to, respectively, the tree-level values of the aforementioned matrix elements of $K(J)$-current operator as a result of the $\mathrm{RI/MOM}$ renormalization condition (3).
Extra care is required in the calculation of the dimensionally-regularized bare matrix elements and the subsequent extraction of renormalization constants entering the r.h.s. of (27) and (28), because of the notorious issue in treating $\gamma_{5}$, and to less extent $\epsilon^{\mu\nu\rho\sigma}$ in DR, especially so in the present case due to the involvement of axial anomaly. In view of the applications related to the study of polarized structure functions of the proton, we choose a (modified) $\overline{\mathrm{MS}}$ prescription where the gauge invariance and Adler-Bell-Jackiw equation are preserved (hence without chiral symmetry). So far, there is no practical prescription known to the authors to apply consistently an anticommuting $\gamma_{5}$ in DR for computing loop amplitudes with (closed) fermion chains with an odd number of $\gamma_{5}$, especially when the amplitudes are anomalous Chen:2023lus; Chen:2024hlv. On the other hand, the so-called Larin’s prescription Larin:1993tq of $\gamma_{5}$, where $\gamma_{5}$ in DR is treated non-anticommuting tHooft:1972tcz; Breitenlohner:1977hr; Breitenlohner:1975hg; Breitenlohner:1976te while the spacetime-metric tensors resulting from contracting a pair of $\epsilon^{\mu\nu\rho\sigma}$ are all in D-dimensions, is particularly suited for dealing with the QCD corrections at hand: both gauge invariance and the Adler-Bell-Jackiw equation are maintained and it is technically convenient to apply at high loop orders in QCD.
With this set-up, we have derived the perturbative expressions for the bare matrix elements involved in (28) up to three loops, truncated to high orders in $\epsilon$ sufficient for obtaining their renormalized counterparts to $\mathcal{O}(\epsilon^{0})$; they are subsequently employed to extract the perturbative results for the finite $R_{ij}$ according to both (28) and (27), with perfect agreement found between the two. (For more technical details of the computational workflow and evaluation of loop integrals, we kindly refer to refs. Ahmed:2021spj; Chen:2021gxv.) For the sake of references, we document below the perturbative results for the finite matching coefficients $R_{ij}$ up to three loops in Landau gauge.
@@
The dependence of $R_{ij}$ on ${\mu}^{2}$ exhibited in the above perturbative expressions are governed by their RG equations. Following from the definition (27) the RG equations for the matching coefficients $R_{ij}$ read
@@
where the matrix of anomalous dimensions in r.h.s. is fully determined by the (proper or modified) $\overline{\mathrm{MS}}$ renormalization constants in (27) in Larin’s prescription Larin:1993tq; these anomalous dimensions observe the relations Breitenlohner:1983pi; Ahmed:2021spj; Luscher:2021bog $\gamma_{11}=-\frac{\mathrm{d}\ln a_{s}}{\mathrm{d}\ln{\mu}^{2}}=-\beta$ and $\gamma_{22}=-2a_{s}\,n_{f}\,\gamma_{12}$ known now to four Chen:2021gxv and five loops Chen:2022lun
With the explicit expression (29) for $R_{11}$, it is interesting to observe that in the limit of large $\beta_{0}$ or $n_{f}$, the leading terms in powers of $\mathcal{O}(a^{n}_{s}\,\beta^{n}_{0})$ exhibit the following neat pattern:

### v2

@@ Appendix A Supplemental Material
$\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}\equiv\begin{pmatrix}\bar{Z}^{\mathrm{DR}}_{11}&\bar{Z}^{\mathrm{DR}}_{12}\\ \bar{Z}^{\mathrm{DR}}_{21}&\bar{Z}^{\mathrm{DR}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,DR}}_{11}&Z^{\mathrm{RI,\,DR}}_{12}\\ Z^{\mathrm{RI,\,DR}}_{21}&Z^{\mathrm{RI,\,DR}}_{22}\end{pmatrix}^{-1}\Big|_{\epsilon\rightarrow 0}=\begin{pmatrix}\bar{Z}^{\mathrm{lat}.}_{11}&\bar{Z}^{\mathrm{lat}.}_{12}\\ \bar{Z}^{\mathrm{lat}.}_{21}&\bar{Z}^{\mathrm{lat}.}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,lat}}_{11}&Z^{\mathrm{RI,\,lat}}_{12}\\ Z^{\mathrm{RI,\,lat}}_{21}&Z^{\mathrm{RI,\,lat}}_{22}\end{pmatrix}^{-1}\,.$ (18)
Consequently, the $R_{ij}$ determined using the first equality $\bar{Z}^{\mathrm{DR}}\cdot\big(Z^{\mathrm{RI,\,DR}}\big)^{-1}\big|_{\epsilon\rightarrow 0}$ with DR can be employed, together with the input $Z^{\mathrm{RI,\,lat}}_{ij}\equiv Z^{\mathrm{RI}}_{ij}$ computed in Eq. (A.1), to obtain the values for the $\overline{\mathrm{MS}}$-renormalization constants $\bar{Z}^{\mathrm{lat}.}_{ij}$ of the operator $K^{\mu}$ and $J_{5}^{\mu}$ but with UV-divergences regularized by the lattice, which can then be applied directly to the bare operator matrix elements computed on the lattice to obtain the $\overline{\mathrm{MS}}$-renormalized results. By virtue of Eq. (16), we can further derive the following shortcut expressions for $R_{ij}$:
@@
where $\overline{K}(\overline{J})_{gg(\bar{q}q)}$ denote the Lorentz-invariant light-cone form-factors or matrix elements of the $\overline{\mathrm{MS}}$-renormalized $K(J)$-current operator but with $\mathrm{RI/MOM}$-renormalized external $g(q)$-fields. Note that $K^{\mathrm{RI}}_{gg}$ and $J^{\mathrm{RI}}_{\bar{q}q}$ are equal to, respectively, the tree-level values of the aforementioned matrix elements of $K(J)$-current operator as a result of the $\mathrm{RI/MOM}$ renormalization condition Eq. (16).
Extra care is required in the calculation of the dimensionally-regularized bare matrix elements and the subsequent extraction of renormalization constants entering the r.h.s. of Eq.(18) and Eq.(19), because of the notorious issue in treating $\gamma_{5}$, and to a lesser extent $\epsilon^{\mu\nu\rho\sigma}$ in DR, especially in the present case due to the involvement of the axial anomaly. In view of the applications related to the study of polarized structure functions of the proton, we choose a (modified) $\overline{\mathrm{MS}}$ prescription where the gauge invariance and Adler-Bell-Jackiw equation are preserved (hence without chiral symmetry). So far, there is no practical prescription known to the authors to apply consistently an anticommuting $\gamma_{5}$ in DR for computing loop amplitudes with (closed) fermion chains with an odd number of $\gamma_{5}$, especially when the amplitudes are anomalous Chen (2023, 2025). On the other hand, the so-called Larin’s prescription Larin (1993) of $\gamma_{5}$, where $\gamma_{5}$ in DR is treated non-anticommuting ’t Hooft and Veltman (1972); Breitenlohner and Maison (1977a, b, c) while the spacetime-metric tensors resulting from contracting a pair of $\epsilon^{\mu\nu\rho\sigma}$ are all in D-dimensions, is particularly suited for dealing with the QCD corrections at hand: both gauge invariance and the Adler-Bell-Jackiw equation are maintained and it is technically convenient to apply at high loop orders in QCD.
With this set-up, we have derived the perturbative expressions for the bare matrix elements involved in Eq.(19) up to three loops, truncated to high orders in $\epsilon$ sufficient for obtaining their renormalized counterparts to $\mathcal{O}(\epsilon^{0})$; they are subsequently employed to extract the perturbative results for finite $R_{ij}$ according to both Eq.(19) and Eq.(18), with perfect agreement found between the two. (For more technical details of the computational workflow and evaluation of loop integrals, we kindly refer to Refs. Ahmed et al. (2021); Chen and Czakon (2022a).) For the sake of references, we document below the perturbative results for the finite matching coefficients $R_{ij}$ up to three loops in Landau gauge.
@@
The dependence of $R_{ij}$ on ${\mu}^{2}$ exhibited in the above perturbative expressions are governed by their RG equations. Following from the definition Eq. (18) the RG equations for the matching coefficients $R_{ij}$ read
@@
where the matrix of anomalous dimensions on the r.h.s. is fully determined by the (proper or modified) $\overline{\mathrm{MS}}$ renormalization constants in Eq. (18) in Larin’s prescription Larin (1993); these anomalous dimensions observe the relations Breitenlohner et al. (1984); Ahmed et al. (2021); Lüscher and Weisz (2021) $\gamma_{11}=-\frac{\mathrm{d}\ln a_{s}}{\mathrm{d}\ln{\mu}^{2}}=-\beta$ and $\gamma_{22}=-2a_{s}\,n_{f}\,\gamma_{12}$ known now to four Chen and Czakon (2022a) and five loops Chen and Czakon (2022b)
With the explicit expression Eq. (20) for $R_{11}$, it is interesting to observe that in the limit of large $\beta_{0}$ or $n_{f}$, the leading terms in powers of $\mathcal{O}(a^{n}_{s}\,\beta^{n}_{0})$ exhibit the following neat pattern: